### v1

# Junhao Chen* 1122    Xiang Li* 33    Xiaojun Ye 1144    Chao Li 55 Zhaoxin Fan 66

Abstract
In this paper, we pursue a novel 3D AIGC setting: generating 3D content from IDEAs. The definition of an IDEA is the composition of multimodal inputs including text, image, and 3D models. To our knowledge, this challenging and appealing 3D AIGC setting has not been studied before. We propose the novel framework called Idea-2-3D to achieve this goal, which consists of three agents based upon large multimodel models (LMMs) and several existing algorithmic tools for them to invoke. Specifically, these three LMM-based agents are prompted to do the jobs of prompt generation, model selection and feedback reflection. They work in a cycle that involves both mutual collaboration and criticism. Note that this cycle is done in a fully automatic manner, without any human intervention. The framework then outputs a text prompt to generate 3D models that well align with input IDEAs. We show impressive 3D AIGC results that are beyond any previous methods can achieve. For quantitative comparisons, we construct caption-based baselines using a whole bunch of state-of-the-art 3D AIGC models and demonstrate Idea-2-3D out-performs significantly. In 94.2% of cases, Idea-2-3D meets users’ requirements, marking a degree of match between IDEA and 3D models that is 2.3 times higher than baselines. Moreover, in 93.5% of the cases, users agreed that Idea-2-3D was better than baselines. Codes, data and models will made publicly available.
Keywords:
LMM agents 3D AIGC Automated 3D Design
11footnotetext: * Indicates Equal Contribution. Indicates Corresponding Author.
Figure 1: The Idea-2-3D framework synergizes the capabilities of the Large Multimodal Model (LMM), Text-to-Image (T-2-I), and Image-to-3D (I-2-3D) models to transform complex multimodal input (one sentence and two images in this case) IDEAs into tangible 3D models. This process begins with the user articulating high-level 3D design requirements (IDEA). Following this, the LMM generates textual prompts (Prompt Generation) that are then converted into 3D models. These models are evaluated through a Multiview Image Generation and Evaluation process, leading to the Selection of an Optimal 3D Model. Subsequently, the T-2-I prompt is refined (Feedback Generation) using insights from the GPT-4V. Additionally, an integrated memory module (see Sec. 3.7), while not depicted here, meticulously records each iteration, facilitating a multimodal, iterative self-refinement cycle within the framework. Note that this procedure is fully automatic without any human intervention.
1 Introduction
Recently the success of 2D AIGC foundataion models [40] has been translated to the 3D domain [20, 22, 33]. However, state-of-the-art models take RGB images or text prompts as inputs, which still fails to match the (wildest) creativity of humanity. Arguably, the nature of creativity is connecting (seemingly unrelated) dots that share intrinsic harmony. So we propose a novel 3D AIGC setting in which all prior arts fail: generating textured 3D models from IDEAs. The formal definition of an IDEA is an interleaved sequence of multi-modal inputs, covering modalities like text, images and 3D models. We show a typical IDEA in Fig. 1, which contains one sentence and two images. This kind of IDEAs come into our minds now and then in the daily life: Someone takes a doughnut into the room and looks at her pet rabbit, imaging what it may look like if the rabbit is eating the doughnut using front paws. This is the moment that creativity happens but as far as we know, no existing 3D AIGC foundation models can take this kind of IDEAs as input. More different types of IDEAs are shown in Fig.2.
Existing methods in text-based 3D model generation [50, 33, 17, 49, 25, 48, 20, 34, 22], known as Text-to-3D (T-2-3D), have made progress in certain aspects such as fidelity, but they still face substantial challenges, particularly when dealing with complex and abstract interleaved multimodal inputs (IDEAs). A viable solution that adapts existing T-2-3D methods for IDEA inputs is to turn IDEAs into natural language descriptions. However, this solution can be time-consuming and demands a specific level of expertise for human users.
Our proposal is to use a multi-agent collaboration framework. LLM (Large Language Model) agent systems [54, 38, 46, 39] have already demonstrated remarkable effectiveness in solving complex natural language processing tasks, suggesting their potential application in the T-2-3D domain. There are already some recent successful methods that leverage LLM agents for computer vision applications, like [10, 45, 51]. They exploit the generic methodology of prompting LLM agents to write codes and invoke existing computer vision functions. We inherit this methodology but exploit LMMs (Large Multimodal Models) as agents because the visual inputs are critical in understanding IDEAs.
However, designing a LMM agent system to generate 3D models from IDEAs is not straightforward and presents its own set of challenges, especially effective integration and understanding of multimodal inputs. To tackle these challenges, we propose Idea-2-3D, a framework that employs three different agents based upon the powerful LMM GPT-4V(ision) [26], for iterative self-improvement in automated 3D design and generation. Specifically, as shown in Fig. 1, Idea-2-3D consists of three LMM agents (green boxes indexed by 1,5,6) acting in the roles of prompt generation, model selection and feedback reflection.
As inspired by Idea2img [55], Idea-2-3D combines the capabilities of LMM agents based upon GPT-4V and other multimodal algorithmic modules (purple boxes indexed by 2,3 and yellow box indexed by 4 in Fig. 1) to generate textual prompts from interleaved user inputs (IDEAs), which are then converted into 3D models. This process involves iterative refinement, utilizing a memory module to record each iteration and support continuous improvement. As shown in Fig. 2, Idea-2-3D can generate high-quality 3D models that well align input IDEAs in a fully automatic manner while caption-based baselines constructed from prior T-2-3D models can hardly generate meaningful results.
Our qualitative comparisons and quantitative experiments demonstrate the effectiveness of Idea-2-3D, especially in handling complex and challenging IDEA inputs, where users perceive that 94.2% of the cases are satisfying IDEA requirements. This is also more than twice as good as existing methods (including closed-source commercial methods) at satisfying IDEA requirements.
To summarize, we make the following contributions:
  • Addressing the novel task of fully automated 3D AIGC by transforming high-level, abstract user IDEAs into concrete 3D models, for the first time.
  • Going further than existing LLM agent systems, we demonstrating the effectiveness of agents based upon LMMs represented by GPT-4V in improving, evaluating, and validating multimodal content for 3D model generation.
  • Introducing Idea-2-3D, which leverages T-2-3D model tools to handle interleaved multimodal inputs (IDEAs) and generate high-quality 3D models.
  • Presenting an evaluation set of challenging multimodal IDEAs and confirming Idea-2-3D’s effectiveness through a comprehensive user preference study.
Figure 2: Overview of 3D models generated from various types of multimodal IDEA inputs supported by Idea-2-3D. The light red box on the left is the user input IDEA containing text, images and 3D models. In the center are the baseline results generated directly from the same T-2-I model with caption-based T-2-I prompt (see  LABEL:Compared_Model_Variants). The model on the right is the result generated by iteratively self-refining the T-2-I prompts with Idea-2-3D. Comparison with more existing methods is shown in  LABEL:fig:comp.
2 Related Works
Self-refining Agents. Our research builds on the self-refinement capability of large language models (LLMs). Recent studies show that LLMs are effective at self-refinement, when exploring unknown environments and tasks [54, 24, 32, 43]. For example, projects such as Self-refine [24] and Reflexion [43] utilize LLMs to iteratively critique their outputs and use feedback to improve prediction.
However, these approaches mainly excel in tasks dealing with natural language descriptions [44]. In contrast, our Idea-2-3D project employs an iterative self-refinement system based on LMM in a multimodal environment, especially for interleaved inputs of text, images, and 3D models (IDEAs), in a different way other than the traditional approach focusing solely on natural language inputs.
Large Multimodal Model. Building on the Large Language Model (LLM), the development of the Large Multimodal Model (LMM) marks an important evolution from unimodal to multimodal processing capabilities. Initial LLMs, such as the GPT family, focused on the generation of textual data, demonstrating superior capabilities in understanding and creating natural language [3, 36].
As the field evolves, the processing capabilities of the models expanded from pure text to include multimodal data including images, audio, and video [37, 35]. For example, CLIP [35] demonstrates cross-modal understanding between text and image, and LLaVA [19] demonstrates multimodal understanding and generation of mixed text, image, and video inputs. Further, projects such as Uni-3D [57] and SDFusion [5] extend this concept to the 3D design domain, enabling good 3D model understanding, generation and reconstruction. LMM has also enabled open-set scene understanding in various settings [47, 15, 16, 21, 11].
Despite the progress made in the field of 3D understanding and generation, current LMMs can still only handle input content at the text and image level, and struggle to handle multimodal high-level inputs containing text, images, and 3D models (IDEAs in our case). In contrast, the Idea-2-3D project takes a much larger step forward in multimodal 3D model generation. Our system is capable of processing not only single-modal inputs, but also composite multimodal inputs containing text, images, and 3D models at the same time.
Extensions of T-2-3D Models. There is already a large literature extending seminal T-2-3D models, including variants enabling T-2-3D models to better follow user prompts [1, 4, 6] , refine keywords in T-2-3D prompts for better visual quality [8], support for additional image inputs for image processing [2, 12], support for additional 3D model inputs for 3D model processing [5], 3D style migration [31, 23, 41], visual concept customization [52, 13], and more. Going even further, Idea-2-3D offers users a more natural way to design and create the 3D content they want. Similar to Visual ChatGPT [53] , which extends ChatGPT’s [30] ability to understand and generate 2D images, Idea-2-3D is designed to provide a more unified and broadly applicable framework for automated 3D model design and generation. Idea-2-3D extends the multimodal input and 3D model understanding and generation capabilities of the LMM [26] using the T-2-I model [42] and I-2-3D model [37], as well as the multimodal input capabilities of the T-2-3D model [50, 33, 17, 49, 25, 48, 34, 20, 22] and other standard off-the-shelf algorithmic modules like multiview rendering.
Figure 3: Overview of the framework of Idea-2-3D, which employs LMM agents to unleash the T-2-3D model’s potential through iterative self-refinement to provide better T-2-3D prompts for the input user IDEA. Green rounded rectangles indicate steps completed by GPT-4V [26] (a.k.a. LMM agents). Purple rounded rectangles indicate T-2-3D modules, including T-2-I models and I-2-3D models. The yellow rounded rectangle indicates the off-the-shelf 3D model multi-view generation algorithm. The blue color indicates the memory module, which saves all the feedback from previous rounds, the best 3D model, and the best text prompt. Note that this cycle is fully automatically executed by LMM agents, without any human intervention.
3 Idea-2-3D Framework
The Idea-2-3D framework represents a novel approach to generate detailed 3D models from high-level, abstract multimodal inputs (IDEAs), shown in Fig. 3. It integrates three LMM-based agents and several off-the-shelf tools for agents to invoke. Specifically, three agents are responsible for prompt generation (green box indexed by 1 in Fig. 3), model selection (green box indexed by 5 in Fig. 3) and feedback generation (green box indexed by 6 in Fig. 3). Two foundation models for T-2-I (purple box indexed by 2 in Fig. 3) and I-2-3D (purple box indexed by 3 in Fig. 3) are exploited together to turn natural language prompts into textured 3D models. As shown in Table. 1, various foundation model variants are evaluated for comprehensiveness. A unique memory module enhances the system, retaining insights from previous iterations to optimize future outputs.
The process begins with the LMM (agent 1) converting multimodal IDEAs into T-2-I prompts, which facilitate the creation of preliminary 3D drafts. These drafts undergo a selection process (agent 5) where the best 3D model is either considered finalized or subjected to further refinement based on LMM feedback (agent 6). The cycle continues until the model satisfies the user’s IDEA (as judged by agent 5) or reaches a pre-set iteration limit. This framework, illustrated in our system diagram (Fig. 3), enables continuous improvement through a loop with automatically generated feedbacks.
3.1 Multimodal IDEA Input
The user-provided IDEA $X$ encapsulates the overarching 3D modeling requirement, represented by a multimodal input set composed of text, images, and 3D models. $X=\{T,I,M\}$, where $T$ is a suite of textual directives encompassing descriptive phrases, keywords, and design specifications that may refer to both 2D and 3D information. $I$ is an assortment of images such as reference shots, diagrams, or related illustrations. $M$ is a compilation of 3D models, including pre-existing constructions or particular design ingredients furnished by the user. Each element of $X$ reflects a facet of the user’s design intent, and their aggregation forms a comprehensive IDEA. This input specification is designed to capture the user’s intent in a multi-faceted and multimodal manner, laying the ground for the subsequent procedural stages, including initial prompt formulation and 3D model synthesis. As mentioned before, designing from this kind of abstract IDEAs is of great needs and our method is the first to fulfill this need.
3.2 Initial Prompt Generation
[…]
$I^{\prime}={\rm CM2I}(M)$ (1)
[…]
$X^{\prime}=\{T,I^{\prime}\cup I\}$ (2)
The LMM agent 1 for prompt generation receives the IDEA $X^{\prime}$ and generates $N$ specific descriptive instructions $\{P_{0},P_{1},...,P_{N-1}\}$:
$P=f_{\rm LMM}(X^{\prime},p_{\rm gen})$ (3)
Here, $p_{\rm gen}$ represents a prompt facilitating the generation of T-2-3D prompts [^2] Note here $p_{\rm gen}$ and P are prompts for LMMs and T-2-3D models, respectively.. This enables Idea-2-3D to extract and interpret not only direct descriptions (e.g. $X^{\prime}$ with only $T$ is also acceptable) but also high-level concepts and intermixed modalities within the IDEA, such as images and 3D models.
In subsequent iterations as denoted by $\rm iter$, each T-2-3D prompt in the set of $P^{\rm iter}$ is used as input to the T-2-3D model to produce draft 3D models $D_{i}^{\rm iter}=T23D(P_{i}^{\rm iter})$, iteratively refining until the output aligns with the user’s intents.
3.3 3D Model Generation
Idea-2-3D transforms the set of N text prompts $\{P_{0},P_{1},\ldots,P_{N-1}\}$ into an equivalent set of N 3D models $\{D_{0},D_{1},\ldots,D_{N-1}\}$ utilizing T-2-3D models. As outlined in Sec. 3.2, T-2-3D encompasses a two-step conversion process: initial Text-to-Image (T-2-I) generation followed by Image-to-3D (I-2-3D) generation. To improve 3D model creation quality, we employ a background removal module [7] on T-2-I outputs before I-2-3D processing. The T-2-3D function is :
$T23D(P_{i})=I23D\circ\text{rembg}\circ T2I(P_{i})$ (4)
In detail: (1) Text-to-Image model $T2I(*)$: Each prompt $P_{i}$ generates a 2D image $G_{i}=T2I(P_{i})$. (2) Background Removal $\rm rembg(*)$: The generated image $G_{i}$ undergoes background removal $G^{\prime}_{i}={\rm rembg}(G_{i})$, enhancing the focus on foreground (i.e., the primary subject). (3) Image-to-3D model $I23D(*)$: The refined image $G^{\prime}_{i}$ is input into I-2-3D, producing the 3D model $D_{i}=I23D(G^{\prime}_{i})$. This methodology from text prompt to 3D model via an intermediary image phase, particularly with background removal, ensures a more accurate and intent-aligned 3D reconstruction, elevating the overall quality of the generated models. These functions are tools invoked by LMM agent 1.
3.4 Draft 3D Model Selection
Then the LMM agent 5 for model selection in Idea-2-3D selects the superior draft 3D model $D_{\rm best}$ from the generated set $\{D_{0},D_{1},...,D_{N-1}\}$ based on the fidelity and relevance to the user’s IDEA. This critical step filters out subpar models, ensuring high-quality iterative generations.
$D_{\text{best}}=f_{\text{select}}(D_{i},X^{\prime},p_{\text{select}})$ (5)
Here, $p_{\rm select}$ is the prompt for the Large Multimodal Model (LMM), guiding the selection of the best draft 3D model. This mechanism compensates for the discrepancy observed when high-quality T-2-I outputs do not necessarily translate into satisfactory I-2-3D models. By assessing the semantic coherence and visual quality of N similar draft models, the LMM identifies the best one. This comparative analysis, akin to a find the difference task, is impossible to achieve by conventional techniques, yet state-of-the-art LMMs like GPV-4V [29, 26] have demonstrated reliable performance in this selection process.
3.5 Feedback Generation
After identifying the best draft model $D_{\rm best}$, the LMM agent 5 decides on whether to finalize this model as the result $D^{*}$ or proceed with refinement. In the latter case, the goal is to generate textual feedback $F$ to guide enhancements for $D_{\rm best}$. This decision hinges on whether the iteration count exceeds a maximum threshold $T$ or if the agent believes no further modifications are needed.
$F=f_{\text{LMM}}(D_{\text{best}},X^{\prime},m,p_{\text{fb}})$ (6)
[…]
3.6 Revised Prompt Generation
In the final stage of each iteration (noted as $\rm iter$), the LMM agent 1 comes to the stage again for Revised Prompt Generation. It uses textual feedback $F$ and the memory module $m$ to create $N$ refined 3D model generation prompts $\{P_{0}^{\rm iter+1},P_{1}^{\rm iter+1},...,P_{n-1}^{\rm iter+1}\}$. This step aims to enhance 3D models generated in the next iteration.
$P_{i}^{\rm iter+1}=f_{\text{LMM}}(F,m,X^{\prime},p_{\text{gen}})$ (7)
Here, $p_{\rm gen}$ is the LMM prompt for I-2-3D prompt generation, which is the same with Eq. 3, despite the inputs are augmented with $F$ and $m$. Note an LMM agent can readily handle different inputs by prompting it that there are two different cases: the initialization case without feedback and memory and the refinement case with feeback and memory. Agent 1 leverages the information stored in $m$ and the previous iteration’s feedback $F$ to generate improved prompts that effectively address the issues identified in $F$. For instance, if feedback $F$ indicates specific visual inaccuracies in the best-to-date model, the revised prompts will focus on rectifying these details through enhanced descriptions.
3.7 Memory Module
The memory module $m$ is integral to the Idea-2-3D framework, serving as a repository for data accrued over the iterative process, as shown by the blue rectangle in Fig. 3. It stores feedback, selected draft 3D models, and corresponding text prompts in a structured 3D model-image-text sequence, enabling the LMM to leverage past experiences and insights gained during previous iterations.
$m_{t}=\{P_{0}^{*},D_{0}^{*},F_{0},\ldots,P_{t-1}^{*},D_{t-1}^{*},F_{t-1}\}$ (8)
In this representation, $P_{\rm iter}^{*},D_{\rm iter}^{*},F_{\rm iter}$ represent the optimal text prompt, 3D model, and textual feedback from each $\rm iter^{th}$ iteration. The memory module $m$ aids the agent in identifying specific T-2-3D output traits, such as misunderstood keywords. This knowledge is then integrated into generating refined 3D model prompts, enhancing the precision and adaptability of model generation. If T-2-3D struggles with certain design aspects, $m$ guides subsequent iterations to optimize prompts more effectively, ensuring continuous improvement in Idea-2-3D, thereby increasing its alignment with complex IDEAs.
4 Experiments
4.1 Experimental Setup
Now we describe our experiment settings. Recall that our method exploits LMM-based agents and off-the-shelf T-2-I and I-2-3D models. As shown in Table. 1, our LMM agent is by default based upon GPT4-V if not specified but we also present results using LLaVA since GPT4-V is not open-sourced. We also experiment with different T-2-I and I-2-3D models to provide as many data points as possible.
T-2-3D baseline. Our first baseline is caption-based. Since no former 3D AIGC methods can be used for IDEA input, we convert image inputs and 3D model inputs (after multiview rendering) into textual descriptions by captioning them with the LMM. In Fig. 2 and Table. 1, this baseline is called T-2-3D.
Ours w/o iterative refinement. To demonstrate the impact of an iterative self-refinement design, we construct an Idea-2-3D variant with only one iteration. Compared with caption-based baselines, this one features multiple prompt generation and best 3D model selection. This baseline is noted in Table. 1 as Idea-2-3D in the Initial-round column.
Lastly, the full Idea-2-3D exploits by a default cap of 5 iterations if not specified, noted in Table. 1 as Idea-2-3D in the Iterative self-refined column.
4.2 Evaluation Protocol
Following the evaluation practice of Parti [56], we provide a set of prompts for evaluation. Our prompt set involves 20 diverse user-generated IDEAs, including examples from Fig. 1 and Fig. 2. These IDEAs span a range of complexities: nine were text-only, five featured a single image or 3D model, and six included two images or 3D models as multimodal inputs. Each test case was meticulously designed to represent real-world scenarios.
The primary metric for our evaluation was a user preference study. Participants were presented with various IDEAs alongside multiple 3D model outputs from both caption-based T-2-3D baselines and Idea-2-3D. To aid decision-making, users viewed rotating video representations of each 3D model. The order of presentation for both the cases and the model outputs was randomized to avoid any sequence bias, ensuring an unbiased assessment of the users’ preferences.
For Table. 1, we evaluate two T-2-I mdoels: DeepFloyd IF [14], and DALL·E [27, 28], two I-2-3D models: Zero123 [9, 20] and Wonder3D [22], and two LMM agent options: GPT-4V [29, 26] and LLaVA-v1.5-13b [19, 18].
Figure 4: Comparison between prompts for "Caption-based T-2-I prompt for initial round", "Idea2-3-D prompt for initial round" and "Iterative self-refinement Idea-2-3D end round" and the comparison between generated 3D models.
Figure 5: An example of the process for a 3D model generated through the Idea-2-3D framework. An initial prompt is first generated based on the input IDEA and $p_{\text{gen}}$, then multiple images are generated from the T-2-I model, and then the corresponding draft 3D model is generated for each image separately. The most appropriate 3D model is selected based on $p_{\text{select}}$ and feedback is generated. The results are output after several rounds of iterations. Green rectangles indicate text prompts for T-2-3D. blue rectangles indicate: modification suggestions given by agent based on draft 3D models and $p_{\text{fb}}$.
4.3 Results
4.3.1 Quantitative user study.
Our evaluation, detailed in Tab. 1, compares caption-based T-2-3D baselines and Idea-2-3D using models specified in each row. We ask users to evaluate which model (the T-2-3D model, the first round results of Idea-2-3D, and the end round results of Idea-2-3D) is more satifying, as well as evaluating each 3D model for compliance with IDEA.
Table 1: This user study’s comparative analysis reveals the remarkable superiority of the Idea-2-3D framework over existing caption-based T-2-3D methods.
Comparison stage: T-2-3D vs. Idea-2-3D (Initial-round) Idea-2-3D (Iterative self-refined)
Model Evaluation Which is better
Satisfying
IDEA
Which is better
Satisfying
IDEA
T-2-I I-2-3D LMM T-2-3D Idea-2-3D T-2-3D Idea-2-3D T-2-3D Idea-2-3D T-2-3D Idea-2-3D
DALL·E Zero123 GPT4-V 25.8% 74.2% 41.5% 78.2% 6.5% 93.5% 41.5% 94.2%
DALL·E Zero123 LLaVA 27.5% 72.5% 70.6% 29.7% 70.3% 82.3%
DALL·E Wonder3D GPT4-V 18.8% 81.2% 47.9% 74.3% 3.5% 96.5% 47.9% 91.1%
DALL·E Wonder3D LLaVA 16.3% 83.7% 69.4% 10.3% 89.7% 76.2%
DeepFloyd IF Zero123 GPT4-V 20.0% 80.0% 38.5% 64.1% 28.3% 71.7% 38.5% 73.7%
DeepFloyd IF Zero123 LLaVA 29.0% 71.0% 57.6% 34.0% 66.0% 66.8%
DeepFloyd IF Wonder3D GPT4-V 29.3% 70.7% 32.6% 66.1% 14.8% 85.2% 32.6% 75.9%
DeepFloyd IF Wonder3D LLaVA 26.0% 74.0% 55.3% 18.3% 81.7% 64.5%
The results show that Idea-2-3D markedly enhances user preference scores across a diverse range of T-2-I, I-2-3D, and LMM models. Notably, the Idea-2-3D framework’s initial prompting stage significantly outperforms the caption-based T-2-3D results by effectively decomposing and interpreting the user’s multimodal IDEA, thereby selecting the most suitable 3D model. This improvement is further amplified in the iterative self-refinement stage of Idea-2-3D.
For instance, in scenarios utilizing DALL-E, Zero123, and GPT-4V, Idea-2-3D models were preferred by users in 74.2% of the cases over T-2-3D in the initial round, demonstrating a higher IDEA satisfaction rate (78.2%). Conversely, T-2-3D models achieved only 41.5% satisfaction. In the iterative refinement comparisons, Idea-2-3D models were favored even more (93.5% preference), achieving a remarkable 94.2% satisfaction rate.
These findings were consistent across various T-2-I, I-2-3D, and LMM configurations. Additionally, we observed that stronger T-2-I models, with enhanced language understanding capabilities, contributed to improved performance in Idea-2-3D, suggesting that our framework may enjoy the development of off-the-shelf models it invokes.
4.3.2 Visualization of the effects of the self-iterative refinement process.
Fig. 4 showcase the evolution and refinement of 3D models at different Idea-2-3D stages and how the caption-based baseline works. Fig. 5 shows the iterative self-improvement process of a case in Idea-2-3D. These visual representations illustrate the framework’s effectiveness in aligning with user IDEAs and the incremental improvements achieved through its iterative process.
4.4 Ablation Study
In this section, we provide ablative user studies to shed light on the effectiveness of critical design choices in Idea-2-3D. Qualitative and quantitative results on shown in Fig. 6 and Table. 2 respectively. Note that in previous experiments, we use a default iteration cap of 5, but here this limit is removed so that the end of iterative refinement is decided solely by whether the LMM agent 5 is satisfied or not. As can be seen in the experiments (sixth iteration and seventh iteration) for the Polemon case in Fig. 6, removing critical modules (including memory, feedback and storing previous models in memory) lead to slow quality convergence while our full Idea-2-3D converges within 3 iterations.
Figure 6: Key module ablation across iterations. Note that the full model reaches satisfactory results (judged by the LMM agent 5) at iteration 3.
The quantitative ablative user study in Tab. 2 shows that: (1) Our proposed memory module can effectively prevent the quality of 3D model generation from divergence (deviating from the user input IDEA). (2) Our proposed LMM feedback agent can effectively improve the speed of quality convergence, i.e., the generated 3D model can well align with user input IDEAs in shorter iteration rounds. (3) If the storage information of the previous model in the memory module is removed, there will be a lower quality upper limit (whether models satisfy IDEAs) after convergence.
Table 2: Ablation study for Idea-2-3D components, collectively obtained at standard iteration rounds ($N=5$) using the 21 test samples presented in this paper.
Satisfying(%) iter 1 iter 2 iter 3 iter 4 iter 5
w/o memory 64.87 60.32 55.34 48.98 41.36
w/o feedback 77.83 80.41 82.67 84.03 85.32
w/o previous Model 78.31 84.58 85.87 85.41 84.49
full model 78.20 89.45 95.23 96.18 96.95
Quality comparisons with commercial models. Qualitative and quantitative comparisons with commercial models are presented in Fig. 7 and Table. 3 respectively, showing that our Idea-2-3D enables a new 3D AIGC capability that existing commercial models fail to demostrate.
Efficiency. Since generation from $N$ prompts can run in parallel, the inference speed is agnostic of $N$. Idea-2-3D generates a result in about 20-30 minutes, since the average iteration round is 2-3 and each iteration takes about 10 minutes. For reference, existing commercial methods typically generate a model in about 5-10 minutes, e.g., 3DTopia and One-2-3-45++ takes about 5 minutes on average, while Luma and Meshy take about 10 minutes. Our method is comparable to them in scale (tens of minutes).
Figure 7: Comparison with commercial T-2-3D models. Prompts are the same as Fig. 2.
Table 3: Quantitative comparison with commercial T-2-3D models.
Method CLIP Score normal-$\uparrow$ Which is best normal-$\uparrow$ Satisfying normal-$\uparrow$
3DTopia 0.2512 2.2% 4.5%
One-2-3-45++ 0.2805 4.3% 24.7%
Luma 0.3068 8.7% 43.2%
Meshy 0.2827 6.5% 32.3%
Ours 0.3611 78.3% 94.2%
5 Conclusion
Idea-2-3D, utilizing an LMM agent collaboration framework, revolutionizes 3D AIGC by automating the creation of models from high-level, interleaved multimodal user inputs (IDEAs). This innovative system excels in integrating text, images, and 3D models, underpinned by a unique iterative process that enhances model coherence and visual alignment with IDEAs. User studies underscore its superiority in satisfaction and comparative quality, marking Idea-2-3D as a significant advancement in 3D AIGC and a benchmark for future design tools.